On January 1, 2018, Ivanhoe Corporation signed a 5-year noncancelable lease for equipment. The terms of the lease called for Ivanhoe to make annual payments of $202000 at the beginning of each year for 5 years beginning on January 1, 2018 with the title passing to Ivanhoe at the end of this period. The equipment has an estimated useful life of 7 years and no salvage value. Ivanhoe uses the straight-line method of depreciation for all of its fixed assets. Ivanhoe accordingly accounts for this lease transaction as a finance lease. The minimum lease payments were determined to have a present value of $842313 at an effective interest rate of 10%. In 2019, Ivanhoe should record interest expense of Answer: (round to the nearest whole dollar)
Answer #1
Period Installment paid Interest expense Lease liability paid carrying value at end
1/1/2018 202000 0 202000 842313-202000= 640313
31/12/2018 or 1/1/2019 202000 640313*10% =64031 202000-64031= 137969 640313-137969= 502344
31/12/2019 or 1/1/2020 202000 502344*10%= 50234 202000-50234= 151766 502344-151766= 350578

Interest expense for year ended 2019 = $ 50234

As per accrual concept ,expense should be recognized in the period in which it is incurred ,so interest accrued for the year should be recognized in the same year whether paid or not .

In 2018 ,after first payment ,interest accrued for the year (1/1/2018 -31/12/2018) should be recognized in 2018 irrespective of same will be paid on 1/1/2019 .

same for year 2019.